OpenStudy (anonymous):

Divide and Simplify 4/7 divided by 32/5

OpenStudy (tanya123):

1/280

OpenStudy (tanya123):

Simplify[4/7/32/5]

OpenStudy (anonymous):

Could you explain how you got the answer if you don't mind

OpenStudy (tanya123):

The answer is 1/280. And this is how you Simplify [4/7/32/5]

OpenStudy (anonymous):

0.00357142857 this is what I am coming up with where am I going wrong

OpenStudy (tanya123):

Hold on, @iPwnBunnies

OpenStudy (ipwnbunnies):

\[\frac{4}{7} \div \frac{32}{5}\] = \[\frac{4}{7} * \frac{5}{32}\]

OpenStudy (anonymous):

ok

OpenStudy (ipwnbunnies):

Another tip. You can reduce the 32 in the denominator. There is a factor of 4 in the numerator. 4 goes into 32 eight times.

OpenStudy (anonymous):

Still confused because when I try to figure it I come up with 3 23/35 maybe i'm making this too complicated for myself